The values of electronegativity of atoms A and B are 4 and 1.20 respectively.the percentage ionic character of A-B bond is

Grade:12

1 Answers

Arun
5 years ago
% ionic character = 16( xA - xB) + 3.5 (xA - xB)^2
Where xA and xB are electronegativities of atom A and B respectively. 
% ionic character = 16( 4-1.2) + 3.5(4-1.2)2
% ionic character = 44.8 + 27.44 = 72.24%